(a) Does the function f(x) = x 3 satisfy the hypothesis of the Mean Value Theorem on the interval [−10, 10]? Justify your answer. (b) If so, determine the point c which is quaranteed by the Mean Value Theorem. 


1
Expert's answer
2021-04-20T06:49:59-0400

Given function, "f(x)=x^3"


(a) As the given function is a polynomial function, so It is differentiable and continuous, so by the mean value theorem, there exists a point in the interval [-10,10].


(b) The point c is given by


"f'(c)=\dfrac{f(b)-f(a)}{b-a}=\dfrac{f(10)-f(-10)}{10-(-10)}=\dfrac{(10)^3-(-10)^3}{20}=\dfrac{2000}{20}=100"


"3c^2=100\Rightarrow c=\sqrt{\dfrac{100}{3}}=\dfrac{10\sqrt{3}}{3}"